OpenStudy (anonymous):

how do you solve the inverse (4 + 5x) / (2x-1)

OpenStudy (anonymous):

put (4+5x)/(2x-1)=y (4+5x)=2xy-y 5x-2xy= -4-y. x=-(4+y)/(5-2y). now replace y by x. then inverce of f(x)=(x+4)/(2x-5)

OpenStudy (anonymous):

thank you